# Generalized Distribution Prediction for Asset Returns

### v1

@@ 3.2.3 Loss Function
Following the methodology of [2], we use a custom loss function based on quantile regression to train the model. The original loss function is designed to minimize the error between the true returns and the predicted quantiles. It is defined as
$L_{\mathcal{T}}=\frac{1}{BK}\sum_{\tau\in\mathcal{T}}\sum_{i=1}^{B}\left[\rho_{\tau}\left(r_{i}-\hat{Q}_{r_{i}}(\tau)\right)+\rho_{\tau}\left(\tilde{r}_{i}-\hat{Q}_{\tilde{r}_{i}}(\tau)\right)\right]$ (1)
where $\mathcal{T}$ is the set of quantiles used for the prediction, $B$ is the batch size (i.e., the number of assets in each batch), $K$ is the number of quantiles in the set $\mathcal{T}$, $r_{i}$ is the raw return of the asset, $\tilde{r}_{i}$ is the standardized (volatility-normalized) return of the asset and $\hat{Q}_{r_{i}}(\tau)$ and $\hat{Q}_{\tilde{r}_{i}}(\tau)$ are the predicted quantiles of the raw and standardized returns for quantile $\tau$, respectively Furhtermore, $\rho_{\tau}(\xi)$ is the quantile loss function
@@
where $\xi$ represents the error (residual) between the true return value and the predicted quantile.
@@ 3.2.4 Modified Loss Function
Our proposed model generalizes the model in [2] by generating multiple predictions over a look-ahead window, and we therefore modify the loss function in 1 to account for each future time step $t\in\{1,2,\dots,T\}$ within the prediction horizon. The modified loss function is defined as
$L_{\mathcal{T}}=\frac{1}{BTK}\sum_{\tau\in\mathcal{T}}\sum_{t=1}^{T}\sum_{i=1}^{B}\left[\rho_{\tau}\left(r_{i,t}-\hat{Q}_{\vec{r}_{i,t}}(\tau)\right)+\rho_{\tau}\left(\tilde{r}_{i,t}-\hat{Q}_{\vec{\tilde{r}}_{i,t}}(\tau)\right)\right]$
where $\mathcal{T}$ is the number of future time steps (i.e., the prediction horizon) in the look-ahead window, $r_{i,t}$ and $\tilde{r}_{i,t}$ are the raw and standardized returns of the asset at future time step $t$, and $\hat{Q}_{\vec{r}_{i,t}}(\tau)$ and $\hat{Q}_{\vec{\tilde{r}}_{i,t}}(\tau)$ are the predicted quantiles of the raw and standardized returns for asset $i$ at future time step $t$.
This modification allows us to compute the loss over the entire prediction window, training the model to minimize the quantile loss for each future time step within the horizon. The inclusion of both raw and standardized returns helps ensure that the model learns both absolute and volatility-adjusted relationships.
@@ 3.3.1 Asset Selection
Several asset classes are chosen to create a diverse representation to emphasize the generalizability of the model. An overview can be found in table 1. Four asset classes are chosen: stocks, currencies, commodities, and cryptocurrencies to capture a diverse set of asset classes to test the robustness of the model. Within stocks there are three sub-asset classes for the index the stocks are listed on, S&P 500, Euro Stoxx 50, and Nikkei 225. Full details on the selected assets as well as summary statistics, can be found in Appendix A.
@@ 3.3.2 Feature Engineering
Additional features are computed for each asset to enhance the model’s ability to capture various market dynamics. Table 2 shows an overview of the engineered features. These features span returns, volatility metrics, moving averages, technical indicators, risk metrics, and the asset class to which the asset belongs. These features are common practice in technical analysis and are widely used across studies [7, 10, 6, 12].
@@ 3.3.3 Market Data
In addition to the assets, general market data is incorporated to capture broader trends and enhance the model’s ability to account for macroeconomic factors. An overview of the market data sources can be found in Table 8. This data includes stock indices, commodities, other indices, and treasury yields. For more details on market data, see Appendix B.
@@
3.3.4 Data Split
The data is split into training, validation, and test sets based on the time periods as follows: Training set: All available data up to the end of 2017. Validation set: Data from 2018 to the end of 2019. Test set: Data from 2019 to 2024, which includes the COVID-19 pandemic and the 2022 bear market. To create varying sequence lengths, each asset’s data is randomly divided into sequences with lengths sampled uniformly from the range $[1,30]$ days.
@@ 3.4 Synthetic Data
We generate 10 synthetic datasets per distribution, each consisting of 1000 samples, to assess the model’s robustness under these simulated conditions. Additionally, we generate $g$ synthetic datasets for the market data, where $g$ represents the number of market features. One feature is designed to have a correlation of approximately 0.7 with the synthetic returns to simulate market co-movement For more details see Appendix E.
@@ 3.4.1 Hyperparameter Optimization
The model’s hyperparameters are optimized using Bayesian optimization, with separate sets of hyperparameters for the asset-specific model (Stage 1) and the market data model (Stage 2). Each set of hyperparameters is sampled independently. The common hyperparameters and the specific hyperparameters for each stage are detailed in Table 4.
Table 4: Hyperparameters for Asset-Specific and Market Data Models
@@
Use layer normalization Boolean: $\{1,0\}$
@@ 3.5 Experimental Setup
Our proposed model -which will be referred to as qLSTM- is evaluated and compared against the original model architecture [2] -which will be called qDense-. It should be noted that the comparison is not entirely fair as the original model was trained under different circumstances and utilized different data. As the qDense model has a fixed forecast window of 22 days, the qLSTM model will only utilize a 22-day-ahead window as well. This means that on the test data, the assets are iterated sequentially taking 22 entries at a time, and then the loss of the 22-day-quantiles of both models is compared using the loss function described in the next section. To act as a baseline for both models, we compare them to a simple linear quantile regression (LQR) model  [^1] The LQR model is trained and tested using the same data as qDense [22].
@@ 3.5.1 Training Procedure
The model is trained using the Adam optimizer with a fixed number of 100 epochs. Early stopping is employed with a patience parameter of 10, meaning that if the validation loss does not improve for 10 consecutive epochs, training is halted. The weights that generated the lowest loss on the validation data are chosen.
3.6 Evaluation Metric
The primary evaluation metric is the quantile loss function, which has been slightly modified to disregard the volatility normalization as well as using only a single return for the evaluation. The quantile loss for a given asset is defined as
@@ 3.6 Evaluation Metric
where $\mathcal{T}$ is the set of quantiles used for prediction, and $K=|\mathcal{T}|$ is the number of quantiles, $T$ is the number of time steps in the test data for the given asset, $r_{t}$ is the raw return of the asset at time step $t$, and $\hat{Q}_{r_{t}}(\tau)$ is the predicted quantile of the raw return of the asset for quantile level $\tau$ at time step $t$. $\rho_{\tau}(\xi)$ is the quantile loss function, defined as:
@@
with $\xi=r_{i,t,j}-\hat{Q}_{r_{i,t,j}}(\tau)$ representing the residual (error) between the true return and the predicted quantile.
@@ 4 Results & Discussion
4.1 Hyper parameter Optimization
Table 5 shows the optimal hyperparameters found in the 100 trials on market data, the final model has a total of $228.390$ trainable parameters. The optimal hyperparameters of the qDense model can be found in Appendix C.
Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
Batch size 256
Learning rate 0.0004
Normalization window 34
Raw LSTM layers 4 1
LSTM units 64 128
Dense layers 1 1
Dense units per layer 128 64
Dropout rate 0.1023
Market activation None
Hidden activation Leaky ReLU
Use layer normalization Yes (1)
L1 regularization 0.0009
L2 regularization 0.0010
4.2 Model Performance Evaluation
Table 6: Comparison of qDense, qLSTM, and LQR across various asset types. Percentage differences are calculated using mean percentage difference.
LQR qDense qLSTM qDense - qLSTM LQR - qLSTM
Commodities 1.4850 0.6226 0.3319 0.2908 (60.93%) 1.1532 (126.94%)
Crypto-currencies 2.5176 1.3605 0.9292 0.4313 (37.68%) 1.5884 (92.17%)
S&P 500 1.4832 0.5969 0.3284 0.2684 (58.02%) 1.1547 (127.48%)
Nikkei 225 1.4805 0.6187 0.3298 0.2889 (60.92%) 1.1507 (127.13%)
Euro Stoxx 50 1.4803 0.5983 0.3326 0.2656 (57.07%) 1.1477 (126.61%)
Currency Pairs 1.2086 0.1597 0.1156 0.0441 (32.05%) 1.0930 (165.09%)
Normal Synthetic 0.2400 0.7003 0.3549 0.3454 (65.46%) -0.1149 (-38.62%)
Log Normal Synthetic 0.4163 0.7075 0.3713 0.3362 (62.33%) 0.0450 (11.43%)
Gamma Synthetic 0.5124 0.7392 0.3805 0.3586 (64.06%) 0.1318 (29.53%)
Uniform Synthetic 0.2188 0.5683 0.2993 0.2691 (62.03%) -0.0805 (-31.06%)
Total Market Mean 1.6092 0.6594 0.3946 0.2649 (50.26%) 1.2146 (121.23%)
Total Synthetic Mean 0.3469 0.6788 0.3515 0.3273 (63.54%) -0.0046 (-1.32%)
Total Mean 1.1043 0.6672 0.3773 0.2898 (55.50%) 0.7269 (98.13%)
The results of the model evaluation across various assets and synthetic datasets, measured in quantile loss, are shown in Table LABEL:tab:final-results. Both the qLSTM and qDense models outperform the Linear Quantile Regression (LQR) model on market data. However, LQR manages to outperform both models on some of the synthetic data, as illustrated in Table LABEL:tab:final-results. The qLSTM model consistently outperforms the qDense model, with an average performance improvement of 50% on real-world datasets, 64% on synthetic datasets, and 55% overall.
Figure 1: The PDFs generated by qLSTM and qDense the first 100 out-of-sample observations of BTC-USD
On real-world data, the qLSTM model demonstrates a significant performance boost. For instance, when applied to S&P 500 stocks, the qLSTM model achieves a quantile loss of 0.3284, compared to 0.5969 for the qDense model and 1.4805 for LQR—representing a 58% and 127% improvement, respectively. Similarly, for cryptocurrencies, the qLSTM model reduces quantile loss by 37% compared to qDense and 92% compared to LQR, showcasing its robustness in handling volatile and unpredictable assets. Figure 1 shows the distribution of quantile losses for qLSTM and qDense over time for the real-world data. The qDense distribution undergoes greater shape variation and appears more symmetrical, while the qLSTM model captures more complexity with less variation in shape, indicating its ability to model intricate distribution patterns. This suggests that qLSTM captures dynamic distributions more effectively, whereas qDense tends to produce distributions closer to normality. This is further illustrated in Figure 2, where a single slice of the distributions shows qDense’s smoother, nearly normal distribution, while qLSTM displays a more intricate and asymmetrical shape, which is better aligned with real data.
Figure 2: The PDFs generated by qLSTM and qDense on a randomly chosen out-of-sample observation of BTC-USD
When evaluated on synthetic data, the qLSTM model shows greater improvement over qDense, but still falls short of LQR on average. qLSTM outperforms LQR on log-normal and gamma distributions, improving quantile loss by 11% and 30%, respectively, demonstrating its ability to capture asymmetrical distributions effectively. However, on the normal and uniform distributions, which are largely symmetrical, LQR outperforms qLSTM by 39% and 31%, respectively. This is somewhat expected, given the nice, symmetric behavior of these distributions. Figure 3 shows the first 100 samples of synthetically generated normally distributed data, where neither distribution changes significantly in shape, with LQR showing no notable variation, as expected for i.i.d. data. This reflects how well the models capture the structure of synthetic data. Although LQR performs better on symmetrical distributions, qLSTM still generates somewhat asymmetrical distributions that approach normality. Figure 4 presents a single slice of the distributions produced by LQR and qLSTM, highlighting qLSTM’s simpler distribution in comparison to real-world data, though still exhibiting slight asymmetry. Interestingly, LQR shows a more complex shape in this case, despite achieving lower quantile loss than qLSTM on this dataset.
Figure 3: The PDFs generated by qLSTM and LQR on the first 100 synthetic normally distributed data
Figure 4: The PDFs generated by qLSTM and LQR on a randomly chosen observation of synthetic normally distributed data

### v2

@@ 3.2.3 Loss Function
Following the methodology of [12], we use a custom loss function based on quantile regression to train the model. The original loss function is designed to minimize the error between the true returns and the estimated quantiles. It is defined as
$\mathcal{L}_{\tau}=\frac{1}{BK}\sum_{\tau\in\mathcal{T}}\sum_{i=1}^{B}\left[\rho_{\tau}\left(r_{i}-\hat{Q}_{r_{i}}(\tau)\right)+\rho_{\tau}\left(\tilde{r}_{i}-\hat{Q}_{\tilde{r}_{i}}(\tau)\right)\right]$ (4)
where $\mathcal{T}$ is the set of quantiles used for the estimation, $B$ is the batch size (i.e., the number of assets in each batch), $K$ is the number of quantiles in the set $\mathcal{T}$, $r_{i}$ is the raw return of the asset, $\tilde{r}_{i}$ is the standardized (volatility-normalized) return of the asset and $\hat{Q}_{r_{i}}(\tau)$ and $\hat{Q}_{\tilde{r}_{i}}(\tau)$ are the estimated quantiles of the raw and standardized returns for quantile $\tau$, respectively Furhtermore, $\rho_{\tau}(\xi)$ is the quantile loss function
@@
where $\xi$ represents the error (residual) between the true return value and the estimated quantile.
@@ 3.2.4 Modified Loss Function
Our proposed model generalizes the model in [12] by generating a singe estimation for log returns over a look-ahead window, and we therefore modify the loss function in Eq.(4) to account for each future time step $t\in\{1,2,\dots,T\}$ within the estimation horizon. The modified loss function is defined as
$\mathcal{L}_{\mathcal{T}}=\frac{1}{BTK}\sum_{\tau\in\mathcal{T}}\sum_{t=1}^{T}\sum_{i=1}^{B}\left[\rho_{\tau}\left(r_{i,t}-\hat{Q}_{\vec{r}_{i}}(\tau)\right)+\rho_{\tau}\left(\tilde{r}_{i,t}-\hat{Q}_{\vec{\tilde{r}}_{i}}(\tau)\right)\right]$
where $T$ is the number of future time steps (i.e., the estimation horizon) in the look-ahead window, $r_{i,t}$ and $\tilde{r}_{i,t}$ are the raw and standardized returns of the asset at future time step $t$, and $\hat{Q}_{\vec{r}_{i,t}}(\tau)$ and $\hat{Q}_{\vec{\tilde{r}}_{i,t}}(\tau)$ are the estimated quantiles of the raw and standardized returns for asset $i$ at future time step $t$.
This modification allows us to compute the loss over the entire estimation window, training the model to minimize the quantile loss for each future time step within the horizon. The inclusion of both raw and standardized returns helps ensure that the model learns both absolute and volatility-adjusted relationships.
@@ 3.3.1 Asset Selection
Several asset classes are chosen to create a diverse representation to emphasize the generalizability of the model. An overview can be found in table 1. Four asset classes are chosen: stocks, currencies, commodities, and cryptocurrencies to capture a diverse set of asset classes to test the robustness of the model. Within stocks there are three sub-asset classes for the index the stocks are listed on, S&P 500, Euro Stoxx 50, and Nikkei 225. Full details on the selected assets as well as summary statistics, can be found in Appendix C.1.
@@ 3.3.2 Feature Engineering
Additional features are computed for each asset to enhance the models’ ability to capture various market dynamics. Table 2 shows an overview of the engineered features. These features span returns, volatility metrics, moving averages, technical indicators, risk metrics, and the asset class to which the asset belongs. These features are common practice in technical analysis and are widely used across studies [27, 28, 29, 30].
@@ 3.3.3 Market Data
In addition to the assets, general market data is incorporated to capture broader trends and enhance the model’s ability to account for macroeconomic factors. An overview of the market data sources can be found in Table 3. This data includes stock indices, commodities, other indices, and treasury yields. For more details on market data, see Appendix C.2.
@@ 3.4 Synthetic Data
We generate 10 synthetic datasets per distribution, each consisting of 1000 samples, to assess the model’s robustness under these simulated conditions. Additionally, we generate $g$ synthetic datasets for the market data, where $g$ represents the number of market features. One feature is designed to have a correlation of approximately 0.7 with the synthetic returns to simulate market co-movement For more details see Appendix C.3.
@@ 3.4.1 Hyperparameter Optimization
The models’ hyperparameters are optimized using Bayesian optimization, with separate sets of hyperparameters for the asset-specific model (Stage 1) and the market data model (Stage 2) [31]. Each set of hyperparameters is sampled independently. The common hyperparameters and the specific hyperparameters for each stage are detailed in Table 4. The optimal hyperparameters found for qLSTM and subsequently qHybrid -because they use the same hyperparameters- can be found in Appendix D.1.
Table 4: Hyperparameters for qLSTM
@@
Use layer
normalization Boolean: $\{1,0\}$
@@ 3.5 Experimental Setup
Our proposed models qLSTM, defined in Eq. (2) and qHybrid, found in Eq. (3) is evaluated and compared against the model architecture from [12] -which will be called qDense. It should be noted that the comparison is not entirely fair as the qDense model was trained under different circumstances and utilized different data. However, to ensure fair comparison as much as we can, in this paper, we always train the models on the same setup. As the qDense model has a fixed forecast window of 22 days, the qLSTM and qHybrid models will only utilize a 22-day-ahead window as well. This means that on the test data, the assets are iterated sequentially taking 22 entries at a time, and then the loss of the 22-day quantiles of both models is compared using the loss function described in the next section. To act as a baseline for all models, we compare them to a simple linear quantile regression (LQR) model  [^1] The LQR model is trained and tested using the same data as qDense [32].
3.5.1 Data Split
The data is split into training, validation, and test sets based on the time periods as follows
  • 1.
    Training set: All available data up to the end of 2017.
  • 2.
    Validation set: Data from 2018 to the end of 2019.
  • 3.
    Test set: Data from 2019 to 2024, which includes the COVID-19 pandemic and the 2022 bear market.
To create varying sequence lengths, each asset’s data is randomly divided into sequences with lengths sampled uniformly from the range $[15,30]$ days.
@@ 3.5.2 Training Procedure
The qLSTM model is trained using the Adam optimizer with a fixed number of 100 epochs. Early stopping is employed with a patience parameter of 10, meaning that if the validation loss does not improve for 10 consecutive epochs, training is halted. The weights that generated the lowest loss on the validation data are chosen. The training details for qDense can be found in appendix D.2.
3.6 Evaluation Metrics
To evaluate the performance of our propsoed modesl, qLSTM and qHybrid, we use several measures, which assess the distribution accuracy in different ways as described below.
3.6.1 Quantile Loss
The primary evaluation metric for comparing qLSTM, qHybrid, with qDense and LQR is the quantile loss function, which has been slightly modified to ignore volatility normalization and to use only one return for the evaluation. The quantile loss for a given asset is defined as
@@ 3.6.1 Quantile Loss
where $\mathcal{T}$ is the set of quantiles used for estimation, and $K=|\mathcal{T}|$ is the number of quantiles, $T$ is the number of time steps in the test data for the given asset, $r_{t}$ is the raw return of the asset at time step $t$, and $\hat{Q}_{r_{t}}(\tau)$ is the estimated quantile of the raw return of the asset for quantile level $\tau$ at time step $t$. $\rho_{\tau}(\xi)$ is the quantile loss function, defined as:
@@
with $\xi=r_{i,t,j}-\hat{Q}_{r_{i,t,j}}(\tau)$ representing the residual (error) between the true return and the estimated quantile.
3.6.2 Distribution Metrics
For the following metrics, we utilize the fact that we estimate the quantiles across multiple observations. We used a forecast window of 30 to get the best estimate for the empirical distribution.
RMSE for Moments
We further evaluate the qLSTM and qHybrid on root mean squared error (RMSE) in moments estimation and compare them to the assumption that the moments are unchanged. , We evalute them with respect to the first four moments, mean ($\mu$), standard deviation ($\sigma$), kurtosis ($\gamma$), and skew ($\kappa$). The error is computes using the formula
$\text{RMSE}_{j}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\lambda_{i,j}-\hat{\lambda}_{i,j})^{2}}$
where $N$ is the test sample, $\lambda$ is the realized moment, $\hat{\lambda}$ is the estimated moment and $j$ is the moment number.
Wasserstain Distance
To evaluate the alignment between the estimated distribution of qLSTM, qHybrid, and the observed returns, we incorporate the Wasserstein distance as an additional metric and compare it against the normal distribution as a benchmark. The Wasserstein distance measures the cost of transforming the estimated distribution into the realized distribution. For a estimated(PDF $p(x)$, discretized over grid points $\{x_{i}\}_{i=1}^{N}$, and realized returns $\{r_{i}\}_{i=1}^{M}$, the Wasserstein distance is defined as
[…]
where $P$ is the estimated empirical distribution and $Q$ is the empirical distribution of realized returns. In practice, we approximate this by sampling $M$ estimated values from $p(x)$ and matching them to the $M$ realized returns, resulting in
$W_{1}(P,Q)=\frac{1}{M}\sum_{i=1}^{M}\left|x_{i}^{*}-r_{i}\right|,$
where $\{x_{i}^{*}\}_{i=1}^{M}$ are the sampled values from the estimated PDF, and $\{r_{i}\}_{i=1}^{M}$ are the sorted realized returns. Lower Wasserstein distances indicate better alignment between the estimated and realized distributions.
Continuous Ranked Probability Score
The Continuous Ranked Probability Score (CRPS) evaluates the alignment between the estimated CDF and the observed values. It measures the integrated squared difference between the estimated CDF and the empirical CDF of the observations. The CRPS for a single observation is defined as
[…]
where $F(x)$ is the estimated CDF for a given observation, $x_{\text{obs}}$ is the observed value, and $\mathbf{1}(x\geq x_{\text{obs}})$ is the indicator function that equals 1 if $x\geq x_{\text{obs}}$ and 0 otherwise. For discrete predictions, the CRPS can be approximated as
$\text{CRPS}(F,x_{\text{obs}})=\frac{1}{N}\sum_{i=1}^{N}\left(F(x_{i})-\mathbf{1}(x_{i}\geq x_{\text{obs}})\right)^{2},$
where $\{x_{i}\}_{i=1}^{N}$ represents the discretized grid over which $F(x)$ is evaluated. The mean CRPS across all samples in the test set is calculated as:
Mean $\text{Mean CRPS}=\frac{1}{M}\sum_{j=1}^{M}\text{CRPS}(F_{j},x_{\text{obs},j}),$
where $M$ is the total number of observations, $F_{j}$ is the estimated CDF for observation $j$, and $x_{\text{obs},j}$ is the corresponding observed value.
This metric provides a comprehensive evaluation of the probabilistic calibration of the model by comparing the estimated distribution with the observed results. Lower CRPS values indicate better alignment between the estimated and observed distributions.
Deviation from Quantile Level (VaR Distance)
To evaluate the calibration of the qLSTM and qHybrid models to estimate tail risks, we calculate the deviation from the desired quantile level at specific quantiles $\tau$ and compare it against a Gaussian baseline. This metric measures how far the observed violation rate deviates from the expected quantile level, with smaller deviations indicating better calibration.
The Violation Rate (VR) is defined as
$\text{VR}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(r_{i}<\hat{Q}_{r}(\tau)),$
where $N$ is the total number of observations, $r_{i}$ is the observed return at time $i$,$\hat{Q}_{r}(\tau)$ is the edstimated quantile (VaR) at level $\tau$, and $\mathbf{1}(x<y)$ is the indicator function that equals 1 if $x<y$, and 0 otherwise.
The deviation from the quantile level $\tau$ is then calculated as
$\text{Deviation}=\left|\text{VR}-\tau\right|,$
where VR is the Violation Rate, and $\tau$ is the expected quantile level (e.g., 0.05, 0.01, 0.00075). This deviation is computed for both the estimated quantiles and a Gaussian baseline.
The resulting metric provides a measure of the model’s calibration quality. A perfectly calibrated model would have $\text{VR}=\tau$, leading to a deviation of 0. Larger deviations indicate underestimation or overestimation of tail risk.
@@ 4 Results & Discussion
In this section we present the performance of our proposed models and compare them to baseline models using various metrics.
4.1 Quantile Loss Performance
Table 5: Quantiles loss for synthetic and non-synthetic data across models.
Quantile loss
Data Type Asset Class LQM qDense qLSTM qHybrid
Real World Commodities 1.4815 0.4136 0.3223 0.3693
Cryptocurrencies 2.4931 2.2843 1.0222 1.2323
S&P 500 1.4811 0.4255 0.3150 0.3572
Nikkei 225 1.4739 0.3991 0.3141 0.3552
Euro Stoxx 50 1.4772 0.4184 0.3124 0.3520
Currency Pairs 1.2003 0.0788 0.1528 0.0973
Total Real World 1.6012 0.6699 0.4065 0.4606
Synthetic Normal 0.5029 0.9338 0.4201 0.4572
Log Normal 0.2115 0.7229 0.3230 0.1565
Gamma 0.3260 0.8268 0.3642 0.2781
Uniform 0.3366 0.6864 0.2940 0.3030
Total Synthetic 0.3443 0.7925 0.3503 0.2987
Total 1.0984 0.7190 0.3840 0.3958
The results in Table 5 highlight the effectiveness of different models (LQM, qDense, qLSTM, and qHybrid) across real-world and synthetic datasets. Overall, the qLSTM model demonstrates superior performance, achieving the lowest quantile loss in most real-world asset classes, including commodities, cryptocurrencies, and equity indices such as the S&P 500, Nikkei 225, and Euro Stoxx 50. Notably, for cryptocurrencies, LSTM reduces the quantile loss to 1.0222 compared to 2.2843 for the Dense model and 2.4931 for the Linear model, showcasing its ability to effectively capture the volatility and non-linearity inherent in this asset class. However, for currency pairs, the Dense model performs best, achieving a quantile loss of 0.0788, which may indicate that simpler architectures can suffice for less complex financial data.
For synthetic data, the qHybrid model demonstrates an edge, particularly for skewed and heavy-tailed distributions such as Log Normal and Gamma. It achieves the lowest quantile losses of 0.1565 and 0.2781, respectively, suggesting that the combination of Gaussian assumptions with learned quantiles is advantageous in these scenarios. However, for simpler distributions such as Normal and Uniform, LSTM outperforms the qHybrid model, indicating its ability to generalize across different statistical patterns. In terms of overall performance, LSTM achieves the lowest total quantile loss for real-world data (0.4065), while the qHybrid model marginally outperforms it for synthetic datasets (0.2987 vs. 0.3503). When considering all datasets, the LSTM model maintains the lowest overall total quantile loss (0.3840), reflecting its versatility and robustness.
These findings underscore the strengths of the qLSTM models in handling real-world sequential financial data, particularly for volatile asset classes like cryptocurrencies and equity indices. The qHybrid model’s success with synthetic data highlights its potential for capturing diverse statistical properties, which could complement LSTM’s strengths in real-world applications. Overall, the two proposed models are able to learn the quantile distribution of sequential data.
4.2 Distribution Metric Performance
In this section, we explore the results of the distribution metrics for qLSTM and qHybrid against a baseline. Note for the sake of conciseness and relevance we omit the comparison against synthetic data and focus only on real world data.
Table 6: Root mean square error of moments estimation compared against the assumption that moments are constant.
Asset class
Stat-istic Model Crypto FX Comm odities Euro Stoxx 50 S&P 500 Nikkei 225
$\mu$ qLSTM 0.0118 0.0010 0.0034 0.0041 0.0032 0.0034
qHybrid 0.0119 0.0020 0.0035 0.0043 0.0033 0.0035
Constant 0.0166 0.0013 0.0050 0.0059 0.0048 0.0050
$\sigma$ qLSTM 0.0424 0.0287 0.0085 0.0091 0.0096 0.0068
qHybrid 0.0400 0.0422 0.0088 0.0109 0.0112 0.0084
Constant 0.0330 0.0021 0.0077 0.0101 0.0099 0.0074
$\gamma$ qLSTM 1.1129 0.7186 0.7667 0.7770 0.8350 0.7422
qHybrid 1.1133 0.7030 0.7496 0.7585 0.8145 0.7339
Constant 1.6036 0.9410 1.0207 1.0598 1.1718 1.0596
$\kappa$ qLSTM 8.1838 9.2768 8.9404 8.7593 8.7733 8.8889
qHybrid 5.3220 4.4449 4.2748 4.0337 4.1521 4.1221
Constant 4.3678 2.7381 2.8323 2.6574 3.1615 2.7176
Table 6 presents the root mean square error (RMSE) for moment estimation across the models qLSTM, qHybrid, and Constant -the assumption that moments are constant throughout $-t$ to 00 and $1$ to $t$- and asset classes. For the mean $\mu$, the qLSTM model consistently achieves the lowest RMSE in all asset classes, highlighting its superior ability to accurately estimate the central tendency of the returns. While the qHybrid model performs comparably, it slightly trails the qLSTM in most categories. The Constant model demonstrates the highest RMSE, suggesting it struggles to capture the mean effectively, particularly for cryptocurrencies and equity indices.
For the standard deviation $\sigma$, the performance varies. The qLSTM model outperforms others in most equity indices, commodities, and cryptocurrencies, reflecting its ability to capture volatility dynamics effectively. However, the Constant model achieves the lowest RMSE for currency pairs, suggesting that simpler methods may suffice for less complex financial assets. The qHybrid model provides competitive results but generally underperforms compared to qLSTM.
The results of skewness $\gamma$ indicate that the qLSTM and qHybrid models perform closely, with the qHybrid model having a slight edge in most asset classes. This suggests that qHybrid models may have an advantage in capturing asymmetry in return distributions. However, the Constant model has significantly higher RMSE values, indicating poor performance in estimating skewness, especially for cryptocurrencies and equity indices.
Finally, for kurtosis $\kappa$, the Constant model achieves the lowest RMSE values, suggesting it captures tail behavior more effectively in this metric. Both qLSTM and qHybrid models show higher RMSE values for kurtosis, which may indicate that their emphasis on other moments slightly compromises their accuracy in tail risk estimation. Nonetheless, qLSTM still maintains competitive performance, particularly for more volatile assets like cryptocurrencies.
Table 7: Wasserstein distances for different models and asset classes. Lowest values are highlighted in bold.
Asset class
Model Crypto FX Comm odities Euro Stoxx 50 S&P 500 Nikkei 225
qLSTM 0.0260 0.0179 0.0079 0.0077 0.0081 0.0075
qHybrid 0.0213 0.0194 0.0072 0.0079 0.0080 0.0071
Gaussian 0.0221 0.0017 0.0074 0.0081 0.0079 0.0074
Asset class
Model Crypto FX Comm odities Euro Stoxx 50 S&P 500 Nikkei 225
qLSTM 0.4320 0.4477 0.4422 0.4424 0.4422 0.4421
qHybrid 0.4179 0.4464 0.4331 0.4340 0.4338 0.4334
Gaussian 0.3544 0.4793 0.4308 0.4366 0.4369 0.4349
Table 8: CRPS values for different models, with the lowest values highlighted.
Table 7 presents the Wasserstein distances for different models and asset classes, highlighting the alignment between the estimated and realized distributions. The qHybrid model achieves the lowest distances for most asset classes, particularly for cryptocurrencies (0.0213) and commodities (0.0072), showcasing its ability to provide a close match to the empirical distributions. However, for currency pairs (FX), the Gaussian model outperforms others with a distance of 0.0017, likely due to the simpler statistical patterns in this asset class. The qLSTM model remains competitive, delivering consistently low Wasserstein distances across asset classes, indicating its robustness in capturing the underlying return distributions.
Table 8 evaluates the Continuous Ranked Probability Score (CRPS) for the models, which evaluates the probabilistic calibration of the estimated distributions. The qHybrid model performs best for most asset classes, including FX (0.4464), Euro Stoxx 50 (0.4340), S&P 500 (0.4338), and Nikkei 225 (0.4334), demonstrating its ability to provide well-calibrated probabilistic forecasts. For cryptocurrencies, the Gaussian model achieves the lowest CRPS (0.3544), suggesting its suitability for this highly volatile asset class. The qLSTM model offers competitive results, with scores close to the best-performing models, reflecting its strong calibration across diverse asset types. These results highlight the complementary strengths of the qHybrid and qLSTM models, with qHybrid excelling in calibration and qLSTM providing robust distributional alignment.
Asset class
Level Model Crypto FX Comm odities Euro Stoxx 50 S&P 500 Nikkei 225
0.05 qLSTM 0.1967 0.0499 0.0659 0.0539 0.0609 0.0549
qHybrid 0.0777 0.0492 0.0481 0.0485 0.0506 0.0460
Gaussian 0.0505 0.0505 0.0466 0.0537 0.0536 0.0469
0.01 qLSTM 0.1085 0.0100 0.0206 0.0190 0.0208 0.0176
qHybrid 0.0478 0.0100 0.0203 0.0220 0.0215 0.0198
Gaussian 0.0298 0.0287 0.0262 0.0326 0.0322 0.0268
0.00075 qLSTM 0.0353 0.0008 0.0028 0.0031 0.0024 0.0018
qHybrid 0.0176 0.0008 0.0036 0.0056 0.0047 0.0029
Gaussian 0.0139 0.0113 0.0104 0.0158 0.0157 0.0099
Table 9: VaR values at different levels for various models, with the lowest values highlighted.
Table 9 presents the Value at Risk (VaR) results for three models (qLSTM, qHybrid, and Gaussian) across asset classes at different confidence levels (0.05, 0.01, and 0.00075). At the 0.05 level, the Gaussian model achieves the lowest VaR for most asset classes, particularly for cryptocurrencies (0.0505) and commodities (0.0466). However, the qHybrid model demonstrates competitive performance, achieving the lowest VaR for FX (0.0492), Euro Stoxx 50 (0.0485), S&P 500 (0.0506), and Nikkei 225 (0.0460), indicating its strength in tail risk modeling across diverse assets. The qLSTM model performs reasonably well but tends to have higher VaR values, particularly for volatile assets like cryptocurrencies.
At the 0.01 level, the qHybrid model shows robust performance across most asset classes, achieving the lowest VaR for commodities (0.0203) and S&P 500 (0.0215). The qLSTM model outperforms the other models for FX (0.0100), Euro Stoxx 50 (0.0190), and Nikkei 225 (0.0176), demonstrating its ability to capture extreme tail risks in these scenarios effectively. The Gaussian model remains competitive, particularly for cryptocurrencies, achieving a VaR of 0.0298, highlighting its utility in assets with higher volatility.
At the extreme tail (0.00075 level), the qLSTM model consistently achieves the lowest VaR across most asset classes, including FX (0.0008), commodities (0.0028), Euro Stoxx 50 (0.0031), S&P 500 (0.0024), and Nikkei 225 (0.0018). This suggests that the qLSTM is particularly adept at capturing risks at the far ends of the distribution. The qHybrid and Gaussian models are slightly less effective at this level, though the Gaussian model shows competitive performance for cryptocurrencies (0.0139). These results emphasize the qLSTM’s strength in extreme tail risk estimation and the qHybrid model’s reliability across broader confidence levels.